OpenStudy (anonymous):

A large asteroid crashed into a moon of a planet, causing several boulders from the moon to be propelled into space toward the planet. Astronomers were able to measure the speed of one of the projectiles. The distance (in feet) that the projectile traveled each second, starting with the first second, was given by the arithmetic sequence 26, 44, 62, 80, ... . Find the total distance that the projectile traveled in seven seconds.

OpenStudy (anonymous):

I got an answer of 152 ft in the 7 seconds.. is that correct?

OpenStudy (lgbasallote):

so first term is 26 the common difference is 18 \[a_n = a_1 + (n-1)d\] \[a_7 = 26 + (7 - 1)(18)\] \[a_7 = 26 + 6 \times 18\] \[a_7 = 26 + 108\] \doesnt sound like 152 to me

OpenStudy (lgbasallote):

i think you did 7 x 18?

OpenStudy (anonymous):

yes i did beccause it said 7 seconds haha

OpenStudy (lgbasallote):

you have to do 7-1..it's a common mistake :)

OpenStudy (lgbasallote):

i remember because it's looking for the 7th term so there are 7 terms so n=7

OpenStudy (asnaseer):

hang on guys...

OpenStudy (lgbasallote):

huh?

OpenStudy (anonymous):

i cant find that formula in my notes :/

OpenStudy (asnaseer):

are you sure you are answering the right question here? it says find the "total distance" - not the distance travelled in the 7th sec

OpenStudy (lgbasallote):

uhmm oops

OpenStudy (anonymous):

yes its the distance traveled

OpenStudy (asnaseer):

:)

OpenStudy (lgbasallote):

wait...i wasnt finished o.O

OpenStudy (anonymous):

Find the total distance that the projectile traveled in seven seconds.

OpenStudy (asnaseer):

so you need to use the formula for the sum of an arithmetic series

OpenStudy (asnaseer):

I'll let @lgbasallote show you how...

OpenStudy (lgbasallote):

i was just finding the a7 for \[s_n = \frac{n(a_1 + a_7}{2}\] :p

OpenStudy (lgbasallote):

hehe :D

OpenStudy (lgbasallote):

so what would be the 7th term?

OpenStudy (anonymous):

umm is a1 the first term and a7 the last term?

OpenStudy (lgbasallote):

yep

OpenStudy (anonymous):

so, the last term of whats given? cuz what about the "..." that means there are more terms :/

OpenStudy (anonymous):

ok, im just really confused

OpenStudy (lgbasallote):

yes because there are 7 terms but it only showed 4 terms..so they put a dotdotdot

OpenStudy (lgbasallote):

the last term is what we solved for

OpenStudy (lgbasallote):

that a7 thingy that we solved..that's the last term

OpenStudy (lgbasallote):

then you just apply the formula for arithmetic sequence

OpenStudy (lgbasallote):

sum of arithmetic sequence*

OpenStudy (perl):

General formula an = a1 + (n-1)d Here we have an = 26 + (n-1)*18

OpenStudy (perl):

now we want sum of an Sum (an) = 26*n +(n-1)(n)/2*18

OpenStudy (perl):

sum an from n=1..7 26*7 + 6*7*18/2 =560

OpenStudy (anonymous):

... i dont understand that at allllllllll

OpenStudy (lgbasallote):

do you understand what i did to find the 7th term?

OpenStudy (anonymous):

Does 560 sound right?